5. You just purchased a share of SPCC for
$100. You expect to receive a dividend of
​$5 in one year. If you expect the price after the dividend is paid to be ​$110​,
what total return will you have earned over the​ year? What was your dividend​ yield? Your capital gain​ rate?
The total return you will have earned over the year is
_____​%.
​(Round to two decimal​ places.)
Answer #1

Total return=(End value-Beginning value+Dividend)/Beginning value  

=(110-100+5)/100

=15%

Dividend yield=Dividend/Beginning value

=(5/100)

=5%

Capital gains yield=(End value-Beginning value)/Beginning value

=(110-100)/100

=10%